- Status Report
- Feb 4, 2023
NOAA GOES SATOPS Morning Report 24 Oct 2001
No change in the status of the POES spacecraft.
Multiple NOAA-16 STX3 playbacks were executed yesterday (as part of test activities) during passes at
1519Z and 1824Z.
NOAA-16 rev 05603 / W at 1645Z: SAIP data recorded for EUMETSAT on DTR 3A did not lock-up
on the PACS bit
synch or frame synch during the playback, but the CDA advised that after the pass that the data had been
recorded on the IMUX.
POES operations were nominal over the past 24 hours.
No change in the status of the DMSP spacecraft.
DMSP operations were nominal over the past 24 hours.
No change in the status of the GOES spacecraft.
At 2356Z on October 23, a frame break was observed on GOES-8 and attributed to loss of signal
The GOES-8 East/West stationkeeping maneuver is scheduled for October 30 at approximately 1708Z
(with a backup day of
November 1). The GOES-10 East/West stationkeeping maneuver is scheduled for November 6 at
(with a backup day of November 8).
GOES-9 will be removed from ZAP storage mode (to three-axis control) on November 13 for a two
month monitoring period
for orbital maintenance and to evaluate wheel and imaging performance. The GOES-12 testing schedule
will not change.
Some GOES images will be canceled up to October 30 due to sun avoidance and seasonal eclipse
operations. Cancellation of
images for GOES-8 will occur from 0345Z to 0639Z. Cancellation of images for GOES-10 will occur
from 0800Z to 1030Z.
GOES KOZ/Eclipse schedules can be reviewed and/or downloaded at
GOES operations were nominal over the past 24 hours.
Questions or comments to [email protected]